# Efficient Large Language Models: A Survey

### v2

@@ 2.5.2. Mixture of Experts (MoE)
Mixture of Experts (MoE) represents a sparse methodology utilized prominently in large-scale models like LLMs. It operates on the principle of segmenting a designated task into several sub-tasks, and then developing numerous smaller, specialized models, dubbed experts, with each honing in on a distinct sub-task. Subsequently, these experts collaborate to deliver a consolidated output. For pre-traning or fine-tuning, MoE helps to manage a huge number of parameters efficiently, enhancing the model’s capacity and potentially its performance while keeping the computational and memory requirements relatively manageable. For inference, MoE decreases the inference time by not engaging all experts simultaneously, but rather activating only a select few. Additionally, MoE is capable of minimizing communication between devices in model-distributed scenarios by allocating each expert to an individual accelerator; communication is only necessary between the accelerators that host the router and the relevant expert model (Kaddour et al., 2023).
@@
MoE-based LLMs. Several MoE-based LLMs have been proposed. For example, GShard (Lepikhin et al., 2021) is a MoE-based LLM that offers a refined method to articulate a variety of parallel computation frameworks with minor modifications to the existing model code. It also amplifies a multilingual neural machine translation Transformer model with Sparsely-Gated MoE beyond 600 billion parameters through automatic sharding. Switch Transformer (Fedus et al., 2022) brings forth a switch routing algorithm and crafts intuitively enhanced models, lowering communication and computational expenditures. It encompasses up to one trillion parameters, dividing tasks among up to 2,048 experts, thereby illustrating the scalability and efficacy of the MoE framework. Artetxe et al. (2022) scale sparse language models to 1.1T parameters, discerning superior performance up to this scale in language modeling, zero-shot and few-shot learning in comparison to dense models. This suggests that sparse MoE models are a computationally efficient substitute for traditionally employed dense architectures. BASE Layer (Lewis et al., 2021) defines token-to-expert allocation as a linear assignment problem, allowing an optimal assignment where each expert acquires an equal number of tokens. PanGu- (Ren et al., 2023b) is a MoE-based LLM with 1.085T parameters, transitioned from the dense Transformer model to a sparse one with Random Routed Experts (RRE), and effectively trains the model over 329B tokens utilizing Expert Computation and Storage Separation (ECSS). Lastly, Mixtral 8x7B (Jiang et al., 2023b) is a MoE with 46.7B total parameters. By leveraging the advantage of MoE architecture, Mixtral 8x7B outperforms LLaMA-2 70B on most benchmarks such as MMLU, MBPP, and GSM-8K with 6x faster inference by only using 12.9B parameters of the model per token for inference.
@@
Algorithm-Level MoE Optimization. The efficiency of MoE-based LLMs can be improved at the algorithm level. The technique termed Expert Choice (Zhou et al., 2022) allows experts to pick the top-k tokens instead of having tokens choose the top-k experts, implying that each token can be directed to a variable number of experts while each expert maintains a fixed bucket size. This method demonstrates higher performance in the GLUE and SuperGLUE benchmarks, and outperforms the T5 dense model in 7 out of the 11 tasks. StableMoE (Dai et al., 2022) identifies the issue of altering target experts for identical input during training and addresses this by creating two training phases. Initially, it cultivates a balanced routing strategy, which is then distilled into a decoupled lightweight router. In the following phase, this distilled router is used for a fixed token-to-expert assignment, ensuring a stable routing strategy. X-MoE (Chi et al., 2022) notes that earlier routing mechanisms foster token clustering around expert centroids, indicating a tendency toward representation collapse. It proposes to estimate the routing scores between tokens and experts on a low-dimensional hyper-sphere. Lifelong-MoE (Chen et al., 2023j) finds that MoE increases the capacity of the model to adapt to different corpus distributions in online data streams without extra computational cost, simply by incorporating additional expert layers and suitable expert regularization. This facilitates continuous pre-training of a MoE-based LLM on sequential data distributions without losing previous knowledge. Lastly, Flan-MoE (Shen et al., 2023) promotes the amalgamation of MoE and instruction tuning, observing that MoE models gain more from instruction tuning compared to dense models. In particular, Flan-MoE effectively enlarges language models without demanding an increase in computational resources or memory requirements.
@@
System-Level MoE Optimization. Several system-level optimization techniques have been developed to accelerate the training and inference of MoE-based LLMs. For example, FastMoE (He et al., 2021) is a distributed MoE training system built on PyTorch, compatible with common accelerators. This system offers a hierarchical interface that allows both flexible model design and easy adaptation to various applications, such as Transformer-XL and Megatron-LM. FasterMoE (He et al., 2022) introduces a performance model that predicts latency and analyzes end-to-end performance through a roofline-like methodology. Utilizing this model, it presents a dynamic shadowing technique for load balancing, a concurrent fine-grained schedule for operations, and a strategy to alleviate network congestion by adjusting expert selection for model training. DeepSpeed-MoE (Rajbhandari et al., 2022) has designed a Pyramid-Residual MoE (PR-MoE) to enhance both the training and the inference efficiency of the MoE model parameter. PR-MoE is a dense-MoE hybrid that employs residual connections to optimally utilize experts, managing to reduce the parameter size by up to 3x without sacrificing quality or compute requirements. Additionally, it proposes a distilled variant, Mixture-of-Students (MoS), which can trim model size by up to 3.7x while retaining quality. TA-MoE (Chen et al., 2022a) highlights that current MoE dispatch patterns do not fully leverage the underlying heterogeneous network environment and thus introduces a topology-aware routing strategy for large-scale MoE training that dynamically modifies the MoE dispatch pattern based on the network topology, making it outperform FastMoE, FasterMoE, and DeepSpeed-MoE. EdgeMoE (Yi et al., 2023) presents an on-device inference engine tailored for MoE-based LLMs. It optimizes memory and computation for inference by distributing the model across different storage levels. Specifically, non-expert model weights are stored directly on the edge device, while expert weights are kept externally and only loaded into the device’s memory when necessary. Tutel (Hwang et al., 2023) is a scalable stack for MoE with adaptive parallelism and pipelining features to accelerate training and inference. It employs a consistent layout for MoE parameters and input data, supporting switchable parallelism and dynamic pipelining without any mathematical inconsistencies or tensor migration costs, thus enabling free run-time optimization. SmartMoE (Zhai et al., 2023) focuses on distributed training for MoE. In the offline stage, SmartMoE constructs a search space of hybrid parallelism strategies. In the online stage, it incorporates light-weight algorithms to identify the optimal parallel strategy. Lastly, MegaBlocks (Gale et al., 2023) transforms MoE-oriented computation with block-sparse operations and creates block-sparse GPU kernels to optimize MoE computation on hardware. This leads to training time up to 40% faster compared to Tutel and 2.4x faster than dense DNNs trained with Megatron-LM.
2.5.3. Long Context LLMs
@@
In many real-world applications, such as multi-turn conversations and meeting summarization, existing LLMs are often required to comprehend or generate context sequences that are much longer than what they have been pre-trained with and may result in a degragation in accuracy due to the poor memorization for the long context. The most obvious and direct way to address this issue is to fine-tune LLMs with similar long-sequence data, which is time consuming and computation intensive. Recently, various new methods have been developed to enable LLMs to adapt to longer context lengths in a more efficient way, including extrapolation and interpolation, recurrent structure, window segment and sliding structure, and memory-retrieval augmentation.
@@
Extrapolation and Interpolation. Standard positional encoding methods like absolute positional embeddings (APE) (Vaswani et al., 2017), learned positional embeddings (LPE) (Wang et al., 2022b), relative positional embeddings (RPE) (Shaw et al., 2018), relative positional bias (Raffel et al., 2020), and rotary position embeddings (RoPE) (Su et al., 2021) have advanced the integration of positional information in LLMs. For example, LPE has been used by GPT-3 (Brown et al., 2020) and OPT (Zhang et al., 2022d); RPE was used by Gopher (Rae et al., 2021) and Chinchilla (Hoffmann et al., 2022), whereas RoPE was used by LLaMA-1 and GLM-130B. However, it is still challenging to train LLMs on sequences with a limited maximum length while still ensuring them to generalize well on significantly longer sequences during inference. Given that, techniques based on positional extrapolation (Press et al., 2022; Sun et al., 2022a; Chen et al., 2023d) and positional interpolation (Chen et al., 2023h; Peng et al., 2023c; Li et al., 2023e) have been proposed.
@@
Positional extrapolation strategies extend the encoding of positional information beyond what the model has explicitly learned during training. For example, ALiBi (Press et al., 2022) applies attention with linear biases to attain extrapolation for sequences that exceed the maximum length seen during training. Through applying negatively biased attention scores, with a linearly diminishing penalty based on the distance between the pertinent key and query, as opposed to using position embeddings, it can facilitate efficient length extrapolation. Different from ALiBi (Press et al., 2022), xPOS (Sun et al., 2022a) characterizes attention resolution as a marker for extrapolation and utilizes a relative position embedding to enhance attention resolution, thereby improving length extrapolation. However, these techniques have not been implemented in some of the recent LLMs such as GPT-4 (OpenAI, 2023), LLaMA (Touvron et al., 2023a), or LLaMA-2 (Touvron et al., 2023b). CLEX (Chen et al., 2023d) proposes to generalize position embedding scaling with ordinary differential equations to model continuous dynamics over length scaling factors. By doing so, CLEX gets rid of the limitations of existing positional extrapolation scaling methods to enable long-sequence generation.
@@
Positional interpolation strategies, on the other hand, reduce the scale of input position indices and extend the context window sizes, allowing LLMs to maintain their performance over longer text sequences. For example, Chen et al. (2023h) highlight that extending beyond the trained context length might impair the self-attention mechanism. They suggest a method that reduces the position indices through linear interpolation, aligning the maximum position index with the prior context window limit encountered during the pre-training phase. NTK interpolation (bloc97, 2023) modifies the base of the RoPE, effectively changing the rotational velocity of each RoPE dimension. YaRN interpolation (Peng et al., 2023c) uses a ramp function to blend linear and NTK interpolation in varying proportions across dimensions and incorporates a temperature factor to counteract distribution shifts in the attention matrix due to long inputs. FIRE (Li et al., 2023e) proposes a functional relative position encoding using learnable mapping of input positions to biases and progressive interpolation, ensuring bounded input for encoding functions across all sequence lengths to enable length generalization. PoSE (Zhu et al., 2023) proposes positional skip-wise training that smartly simulates long inputs using a fixed context window and design distinct skipping bias terms to manipulate the position indices of each chunk. This strategy reduces memory and time overhead compared with full-length fine-tuning.
@@
Recurrent Structure. LLMs’ ability to manage long sequences can also be enhanced through recurrence structure. For example, Transformer-XL (Dai et al., 2019) presents a segment-level recurrence mechanism and utilizes enhanced relative positional encoding to capture long-term dependencies and address the long-context fragmentation issue. Memformer (Wu et al., 2020) leverages an external dynamic memory for encoding and retrieving past information, achieving linear time and constant memory space complexity for long sequences. It also proposes Memory Replay Back-Propagation (MRBP) to facilitate long-range back-propagation through time with significantly lower memory requirements. -former (Martins et al., 2022) presents a Transformer model augmented with unbounded long-term memory (LTM), employing a continuous space attention framework to balance the quantity of information units accommodated in memory against the granularity of their representations. Recurrent Memory Transformer (RMT) (Bulatov et al., 2022) uses a recurrence mechanism to retain information from the past segment level by incorporating special memory tokens into the input or output sequence, demonstrating superior performance compared to Transformer-XL in long context modeling. Block-Recurrent Transformers (Hutchins et al., 2022) utilize self-attention and cross-attention to execute a recurrent function across a broad set of state vectors and tokens so as to model long sequences through parallel computation. Lastly, Retentive Network (Sun et al., [n.d.]) introduces a multi-scale retention mechanism as an alternative to multi-head attention. By encompassing parallel and chunk-wise recurrent representations, it results in effective scaling, allows for parallel training, and achieves training parallelization and constant inference cost, while offering linear long-sequence memory complexity compared to other Transformer models.
@@
Segmentation and Sliding Window. Segmentation and sliding window techniques tackle the issue of long-context processing by dividing the input data into smaller segments, or applying a moving window to slide through the long sequence. For instance, Mistral (Jiang et al., 2023b) uses sliding window attention to effectively handle sequences of arbitrary length with a reduced inference cost. StreamingLLM (Xiao et al., 2023b) identifies an attention sink phenomenon, noting that retaining the Key-Value of initial tokens significantly restores the performance of window attention. Based on this observation, it suggests an efficient framework via merging window context and the first token, allowing LLMs trained with a finite length attention window, but have the ability to generalize to infinite sequence lengths without any fine-tuning. Parallel Context Windows (PCW) (Ratner et al., 2023) segments a long context into chunks, limiting the attention mechanism to function only within each window, and then redeploys the positional embeddings across these windows. LongNet (Ding et al., 2023a) proposes dilated attention, which exponentially expands the attentive field as the distance increases, enabling the handling of sequence lengths of more than 1 billion tokens. LongNet can be implemented by parallelizing training by partitioning the sequence dimension. SLED (Ivgi et al., 2023) is a straightforward method for handling long sequences that repurposes and capitalizes on well-validated short-text language models for use in LLMs.
@@
Memory-Retrieval Augmentation. Several studies tackle the inference of extremely long text by employing memory-retrieval augmentation strategies. A notable example is the KNN-augmented Transformer (Wu et al., 2022), which extends the attention context size by utilizing k-nearest-neighbor (KNN) lookup to fetch previously similar context embeddings. Landmark Attention (Mohtashami and Jaggi, 2023) employs a landmark token to represent each block of input and trains the attention mechanism to utilize it for choosing relevant blocks. This allows the direct retrieval of blocks through the attention mechanism while maintaining the random access flexibility of the previous context, demonstrating impressive performance on LLaMA-1 for long-context modeling. LongMem (Wang et al., 2023a) proposes a decoupled network architecture with the original backbone LLM as a memory encoder and an adaptive residual side network as a memory retriever and reader, efficiently caching and updating long-term past contexts to prevent knowledge staleness. Unlimiformer (Bertsch et al., 2023) enhances the KNN-augmented Transformer by outputting attention dot-product scores as KNN distances, enabling the indexing of virtually unlimited input sequences. Focused Transformer (FoT) (Tworkowski et al., 2023) highlights that the ratio of relevant keys to irrelevant ones diminishes as the context length increases and proposes an optimized solution through contrastive learning to refine the structure of the key-value space. Lastly, Xu et al. (2023a) discover that an LLM with a 4K context window, when augmented with simple retrieval during generation, can match the performance of a fine-tuned LLM with a 16K context window using positional interpolation (Chen et al., 2023h) on long context tasks, while requiring significantly less computation.
2.5.4. Transformer-Alternate Architectures

### v4

@@ 2.5.2 Mixture of Experts (MoE)
Mixture of Experts (MoE) represents a sparse methodology utilized prominently in large-scale models like LLMs. It operates on the principle of segmenting a designated task into several sub-tasks, and then developing numerous smaller, specialized models, dubbed experts, with each honing in on a distinct sub-task. Subsequently, these experts collaborate to deliver a consolidated output. For pre-traning or fine-tuning, MoE requires developers to manage a huge number of parameters efficiently, enhancing the model’s capacity and potentially its performance while keeping the computational and memory requirements relatively manageable. For inference, MoE decreases the inference time by not engaging all experts simultaneously, but rather activating only a select few. Additionally, MoE is capable of minimizing communication between devices in model-distributed scenarios by allocating each expert to an individual accelerator; communication is only necessary between the accelerators that host the router and the relevant expert model (Kaddour et al., 2023).
@@
MoE-based LLMs. Several MoE-based LLMs have been proposed. For example, GShard (Lepikhin et al., 2021) is a MoE-based LLM that offers a refined method to articulate a variety of parallel computation frameworks with minor modifications to the existing model code. It also amplifies a multilingual neural machine translation Transformer model with Sparsely-Gated MoE beyond 600 billion parameters through automatic sharding. Switch Transformer (Fedus et al., 2022) brings forth a switch routing algorithm and crafts intuitively enhanced models, lowering communication and computational expenditures. It encompasses up to one trillion parameters, dividing tasks among up to 2,048 experts, thereby illustrating the scalability and efficacy of the MoE framework. Artetxe et al. (2022) scale sparse language models to 1.1T parameters, discerning superior performance up to this scale in language modeling, zero-shot and few-shot learning in comparison to dense models. This suggests that sparse MoE models are a computationally efficient substitute for traditionally employed dense architectures. Its largest MoE model outperforms its dense counterpart where the latter requires twice as much computation. BASE Layer (Lewis et al., 2021) defines token-to-expert allocation as a linear assignment problem, allowing an optimal assignment where each expert acquires an equal number of tokens, achieving lower validation perplexity during training relative to Switch Transformer. PanGu- (Ren et al., 2023b) is a MoE-based LLM with 1.085T parameters, transitioned from the dense Transformer model to a sparse one with Random Routed Experts (RRE), and effectively trains the model over 329B tokens utilizing Expert Computation and Storage Separation (ECSS). It outperforms dense model like ERNIE 3.0 Titan Wang et al. (2021) on zero-shot test of Chinese downstream task. Lastly, Mixtral 8x7B (Jiang et al., 2023a) is a MoE with 46.7B total parameters. By leveraging the advantage of MoE architecture, Mixtral 8x7B outperforms LLaMA-2 70B on most benchmarks such as MMLU, MBPP, and GSM-8K with 6x faster inference by only using 12.9B parameters of the model per token for inference.
@@
Algorithm-Level MoE Optimization. The efficiency of MoE-based LLMs can be improved at the algorithm level. Expert Choice (Zhou et al., 2022) allows experts to pick the top-k tokens instead of having tokens choose the top-k experts, implying that each token can be directed to a variable number of experts while each expert maintains a fixed bucket size. This method demonstrates higher performance in the GLUE and SuperGLUE benchmarks, and outperforms T5 dense model in seven out of the 11 tasks. StableMoE (Dai et al., 2022) identifies the issue of altering target experts for identical input during training and addresses this by creating two training phases. Initially, it cultivates a balanced routing strategy, which is then distilled into a decoupled lightweight router. In the following phase, this distilled router is used for a fixed token-to-expert assignment, ensuring a stable routing strategy. StableMoE shows better results than Switch Transformer and BASE Layer with lower validation perplexity on language modeling. X-MoE (Chi et al., 2022) notes that earlier routing mechanisms foster token clustering around expert centroids, indicating a tendency toward representation collapse. It proposes to estimate the routing scores between tokens and experts on a low-dimensional hyper-sphere, showing improvements over Switch Transformer on multilingual multi-task benchmark. Lifelong-MoE (Chen et al., 2023f) observes that MoE increases the capacity of the model to adapt to different corpus distributions in online data streams without extra computational cost, simply by incorporating additional expert layers and suitable expert regularization. This facilitates continuous pre-training of a MoE-based LLM on sequential data distributions without losing previous knowledge. It outperforms other MoE models such as GShard on natural language generation and understanding tasks. Lastly, Shen et al. (2024) observe that compared to dense models, MoE gains more from instruction tuning. Based on this observation, they propose Flan-MoE, which combines MoE and instruction tuning to enlarge language models without increasing demands in memory and compute resources while showing better zero-shot and few-shot performance compared to FLAN-T5 dense model.
@@
System-Level MoE Optimization. The efficiency of MoE-based LLMs can also be improved at the system level. For example, FastMoE (He et al., 2021) is a distributed MoE training system built on PyTorch, compatible with common accelerators. This system offers a hierarchical interface that allows both flexible model design and easy adaptation to various applications, such as Transformer-XL and Megatron-LM. FasterMoE (He et al., 2022a) tries to address the challenges of dynamic load imbalance, inefficient synchronous execution mode, and congested all-to-all communication during MoE training. It first introduces a performance model that predicts latency and analyzes end-to-end performance through a roofline-like methodology. Utilizing this model, it presents a dynamic shadowing technique for load balancing, a concurrent fine-grained schedule for operations, and a strategy to alleviate network congestion by adjusting expert selection for model training. It outperforms FastMoE and achieves 1.37× - 17.87× speedup compared to methods including ZeRO, GShard, and BASE Layer. Lina (Li et al., 2023a) also improves training efficiency and inference time by optimizing communication and load balancing in distributed MoE. Lina first prioritizes all-to-all over allreduce using tensor partitioning and pipelining to improve its bandwidth in training, and then dynamically balances the workload with token-level expert selection pattern in inference. DeepSpeed-MoE (Rajbhandari et al., 2022) has designed a Pyramid-Residual MoE (PR-MoE) architecture to enhance both the training and the inference efficiency of the MoE model parameter. PR-MoE is a dense-MoE hybrid that employs residual connections to optimally utilize experts, managing to reduce the parameter size by up to 3x without sacrificing quality or compute requirements. It serves massive MoE models with up to 4.5x faster and 9x cheaper inference compared to quality-equivalent dense models. DeepSpeed-MoE also proposes a highly optimized MoE inference system which enables efficient scaling of inference workloads on hundreds of GPUs, providing up to 7.3x reduction in inference latency and cost when compared with existing MoE inference solutions. TA-MoE (Chen et al., 2022a) highlights that current MoE dispatch patterns do not fully leverage the underlying heterogeneous network environment and thus introduces a topology-aware routing strategy for large-scale MoE training that dynamically modifies the MoE dispatch pattern based on the network topology, making it outperform FastMoE, FasterMoE, and DeepSpeed-MoE. EdgeMoE (Yi et al., 2023) presents an on-device inference engine tailored for MoE-based LLMs. It optimizes memory and computation for inference by distributing the model across different storage levels. Specifically, non-expert model weights are stored directly on the edge device, while expert weights are kept externally and only loaded into the device’s memory when necessary. Tutel (Hwang et al., 2023) proposes adaptive parallelism and pipelining features to adapt to the dynamic workload. It employs a consistent layout for MoE parameters and input data, supporting switchable parallelism and dynamic pipelining without any mathematical inconsistencies or tensor migration costs, thus enabling free run-time optimization, achieving up to 5.75× speedup for a single MoE layer. SmartMoE (Zhai et al., 2023) focuses on the automatic parallelization for MoE distributed training. In the offline stage, SmartMoE constructs a search space of hybrid parallelism strategies. In the online stage, it incorporates light-weight algorithms to identify the optimal parallel strategy. It achieves up to 1.88× speedup in end-to-end training over FasterMoE on a distributed training setting. Lastly, MegaBlocks (Gale et al., 2023) transforms MoE-oriented computation with block-sparse operations and creates block-sparse GPU kernels to optimize MoE computation on hardware. This leads to training time up to 40% shorter compared to Tutel and 2.4x shorter than dense models trained with Megatron-LM.
2.5.3 Long Context LLMs
@@
In many real-world applications, such as multi-turn conversations and meeting summarization, existing LLMs are often required to comprehend or generate context sequences that are much longer than what they have been pre-trained with and may result in a degradation in accuracy due to the poor memorization for the long context. One direct way to address this issue is to fine-tune LLMs with similar long-sequence data, which, however, is time consuming and computation-intensive. To fill this gap, new methods have been developed to enable LLMs to adapt to longer context lengths in a more efficient way. These methods can be in general grouped into four categories: extrapolation and interpolation, recurrent structure, segmentation and sliding window, and memory-retrieval augmentation.
@@
Positional Extrapolation and Interpolation. Standard positional encoding methods such as absolute positional embeddings (APE) (Vaswani et al., 2017), learned positional embeddings (LPE) (Wang et al., 2022a), relative positional embeddings (RPE) (Shaw et al., 2018), and rotary position embeddings (RoPE) (Su et al., 2023b) have advanced the integration of positional information in LLMs. For example, LPE has been used by GPT-3 and OPT, RPE was used by Gopher (Rae et al., 2022) and Chinchilla (Hoffmann et al., 2022), whereas RoPE was used by LLaMA-1 and GLM-130B. However, it is still challenging to train LLMs on sequences with a limited maximum length while still ensuring them to generalize well on significantly longer sequences during inference. Given that, techniques based on positional extrapolation (Press et al., 2022; Sun et al., 2023c; Chen et al., 2024a) and positional interpolation (Chen et al., 2023d; Peng et al., 2024; Li et al., 2024a) have been proposed.
@@
Positional extrapolation strategies extend the encoding of positional information beyond what the model has explicitly learned during training. For example, ALiBi (Press et al., 2022) applies attention with linear biases to attain extrapolation for sequences that exceed the maximum length seen during training. By applying negatively biased attention scores with a linearly diminishing penalty based on the distance between the pertinent key and query, it facilitates efficient length extrapolation. Different from ALiBi, xPOS (Sun et al., 2023c) characterizes attention resolution as a marker for extrapolation and utilizes a relative position embedding to enhance attention resolution, thereby improving length extrapolation. However, these techniques have not been implemented in some of the recent LLMs such as GPT-4, LLaMA, and LLaMA-2. CLEX (Chen et al., 2024a) proposes to generalize position embedding scaling with ordinary differential equations to model continuous dynamics over length scaling factors. In doing so, CLEX gets rid of the limitations of existing positional extrapolation scaling methods to enable long-sequence generation.
@@
Positional interpolation strategies, on the other hand, reduce the scale of input position indices and extend the context window sizes, allowing LLMs to maintain their performance over longer text sequences. For example, Chen et al. (2023d) observe that extending beyond the trained context length could impair the self-attention mechanism. They propose RoPE-PI to reduce the position indices through linear interpolation, aligning the maximum position index with the prior context window limit encountered during pre-training. NTK interpolation (bloc97, 2023) modifies the base of the RoPE, effectively changing the rotational velocity of each RoPE dimension. YaRN interpolation (Peng et al., 2024) uses a ramp function to blend linear and NTK interpolation in varying proportions across dimensions and incorporates a temperature factor to counteract distribution shifts in the attention matrix caused by long inputs. Experimental results on long-text modeling shows that YaRN outperforms existing RoPE interpolation methods including RoPE-PI and NTK. FIRE (Li et al., 2024a) proposes a functional relative position encoding using learnable mapping of input positions to biases and progressive interpolation, ensuring bounded input for encoding functions across all sequence lengths to enable length generalization. It demonstrates competitive results compared to ALiBi, RoPE, and RoPE-PI on long text benchmarks. Lastly, PoSE (Zhu et al., 2024) proposes positional skip-wise training that simulates long inputs using a fixed context window and designs distinct skipping bias terms to manipulate the position indices of each chunk. This strategy reduces memory and time consumption compared to full-length fine-tuning.
@@
Recurrent Structure. LLMs’ ability to manage long sequences can also be enhanced through recurrence structure. For example, Transformer-XL (Dai et al., 2019) presents a segment-level recurrence mechanism and utilizes enhanced relative positional encoding to capture long-term dependencies and address the long-context fragmentation issue. Memformer (Wu et al., 2022a) leverages an external dynamic memory for encoding and retrieving past information, achieving linear time and constant memory space complexity for long sequences. It also proposes Memory Replay Back-Propagation (MRBP) to facilitate long-range back-propagation through time with significantly lower memory requirements, achieving better results than Transformer-XL on language modeling and image generation benchmarks. -former (Martins et al., 2022) presents a Transformer model augmented with unbounded long-term memory (LTM). It employs a continuous space attention framework to balance the quantity of information units accommodated in memory against the granularity of their representations, showing better results than Transformer-XL on long text sorting and modeling. Recurrent Memory Transformer (RMT) (Bulatov et al., 2022) uses a recurrence mechanism to retain information from the past segment level by incorporating special memory tokens into the input or output sequence, and demonstrates superior performance compared to Transformer-XL in long context modeling. Block-Recurrent Transformer (BRT) (Hutchins et al., 2022) utilizes self-attention and cross-attention to execute a recurrent function across a broad set of state vectors and tokens so as to model long sequences through parallel computation. Lastly, Retentive Network (Sun et al., 2023b) introduces a multi-scale retention mechanism as an alternative to multi-head attention. By leveraging parallel and chunk-wise recurrent representations, it enables effective scaling, achieves training parallelization and constant inference cost, and offers linear long-sequence memory complexity compared to other Transformer models.
@@
Segmentation and Sliding Window. Segmentation and sliding window techniques tackle the issue of long-context processing by dividing the input data into smaller segments, or applying a moving window to slide through the long sequence. For instance, Mistral (Jiang et al., 2023a) uses sliding window attention to handle sequences of arbitrary length with a reduced inference cost. StreamingLLM (Xiao et al., 2024) identifies an attention sink phenomenon, noting that retaining the Key-Value of initial tokens significantly restores the performance of window attention. Based on this observation, it suggests an efficient framework via merging window context and the first token, allowing LLMs trained with a finite length attention window, but have the ability to generalize to infinite sequence lengths without any fine-tuning. Parallel Context Windows (PCW) (Ratner et al., 2023) segments a long context into chunks, limiting the attention mechanism to function only within each window, and then re-deploys the positional embeddings across these windows. LongNet (Ding et al., 2023a) proposes dilated attention, which exponentially expands the attentive field as the distance increases, enabling the handling of sequence lengths of more than one billion tokens. SLED (Ivgi et al., 2023) handles long sequences by partitioning the long text into small chunks and leveraging pretrained short-text language models for encoding and decoding. Different from the methods mentioned above, MemWalker (Chen et al., 2023c) transforms lengthy texts into segmented summaries within a tree structure, leveraging LLMs as an interactive entity for guided reading through iterative prompts. It outperforms traditional methods based on extended context, recurrence, and retrieval. Similar to MemWalker, RAPTOR (Sarthi et al., 2024) utilizes text chunks to construct a recursive tree to enable information integration from extensive documents across various abstraction levels during inference, achieving superior performance in multi-step reasoning.
@@
Memory-Retrieval Augmentation. Lastly, several studies tackle the inference of extremely long text by employing memory-retrieval augmentation strategies. A notable example is the Memorizing Transformer (Wu et al., 2022c), which extends the attention context size by utilizing k-nearest-neighbor (KNN) lookup to fetch previously similar context embeddings. Additionally, Landmark Attention (Mohtashami & Jaggi, 2023) employs a landmark token to represent each block of input and trains the attention mechanism to utilize it for choosing relevant blocks. This allows the direct retrieval of blocks through the attention mechanism while maintaining the random access flexibility of the previous context, demonstrating comparable perplexity as Transformer-XL while reducing FLOPs for long-context modeling. LongMem (Wang et al., 2023e) proposes a decoupled network architecture with the original backbone LLM as a memory encoder and an adaptive residual side network as a memory retriever and reader, efficiently caching and updating long-term past contexts to prevent knowledge staleness. It outperforms Memorizing Transformer on long text modeling and natural language understanding tasks. Unlimiformer (Bertsch et al., 2023) enhances the KNN-augmented Transformer by outputting attention dot-product scores as KNN distances to enable indexing of virtually unlimited input sequences. Experimental results show that Unlimiformer outperforms Memorizing Transformer on long document summarization benchmarks. Tworkowski et al. (2023) observe that the ratio of relevant keys to irrelevant ones diminishes as context length increases. Based on this observation, they propose Focused Transformer (FoT), a contrastive learning-based technique to refine the structure of the Key-Value space. Unlike Memorizing Transformer and Transformer-XL, FoT does not require training on long sequences and shows better performance on both long context and short context tasks. Lastly, Xu et al. (2024a) discover that an LLM with a 4K context window, when augmented with simple retrieval during generation, can match the performance of a fine-tuned LLM with a 16K context window using positional interpolation (Chen et al., 2023d) on long context tasks while requiring significantly less computation.
2.5.4 Transformer-Alternate Architectures